Find how many quarts of 4% butterfat milk and 1% butterfat milk should be mixed to yield 60 quarts of 2% butterfat milk.
Answer by mananth(16949)

percent ----------------quantity
4 ----------------x
1 ----------------60-x
2.00% ----------------60
...
4x+1(60 -x)=2*60
4x+60-x =120
4x-x=120-60
3x=60
/3
x=20 quarts 4%
40 quarts 1%